OpenStudy (bmk614):

A particle with velocity at any time t given by v(t)=2e^(2t) moves in a straight line. How far does the particle travel during the time interval when its velocity increases from 2 to 4. A) 1 B) 2 C) 3 D) e^4 E) e^8-e^4

OpenStudy (peachpi):

Integrate v(t) from 2 to 4

OpenStudy (bmk614):

That is what I did, and I got E, but the answer is A

OpenStudy (peachpi):

I got E as well

OpenStudy (irishboy123):

" when its velocity increases from 2 to 4."

OpenStudy (irishboy123):

ie not \(2 \le t \le 4\)

OpenStudy (peachpi):

I see. so you have to solve for the t

OpenStudy (bmk614):

So what do I do?

OpenStudy (irishboy123):

get \(t_1\) from \(v(t)=2e^{2t} = 2\) get \(t_2\) from \(v(t)=2e^{2t} = 4\)

OpenStudy (bmk614):

I got \[t _{1}=0, t_2=\ln(2)/2\]

OpenStudy (bmk614):

So t_1-t_2 doesnt equal 1

OpenStudy (irishboy123):

\(t_1 = 0\) \(v(t) = 4 \implies \ln \sqrt{2}\) we agree so now do the integration

OpenStudy (bmk614):

wouldnt it be ln(2)/2

OpenStudy (irishboy123):

yes \(ln(2)/2 = \ln \sqrt{2}\) so we have: \(\Large \int\limits_{0}^{\ln \sqrt{2} } ~ 2 e^{2t} ~ dt\) are you happy with that?

OpenStudy (bmk614):

I plugged that into my calculator and got 1... Thank you so much!

OpenStudy (irishboy123):

hope it's right :-)) lol!!

OpenStudy (bmk614):

Well it gave me the right answer, so I think so haha.